# On Adversarial Attacks In Acoustic Drone Localization

### v1

@@ 3.3.3 Perturbation Propagation
Much like the ”clean” sound emitted by the drone, our adversarial perturbation must also be propagated to the drone’s sensor array to be sampled. Our goal is to formulate a universal attack optimization scheme shaping the emitted signal at the source $s_{\mathrm{p}}$ so that the sampled response at the sensor $\mu$, denoted $\sigma_{\mathrm{p}}$, would optimally reduce localization accuracy. In our experiments, we consider both optimization of the emitted perturbation for a sound source located in the center of the room, as well as optimization of the emitted sound jointly with the perturbation speaker’s location $\mathbf{x}_{\mathrm{p}}\in\mathbb{R}^{2}$. For our localization quality measure, we choose Mean-Squared-Error (MSE), following the criterion from (Shor et al., 2024). Our universal attack optimization, in its most general form, is therefore formulated according to Equation 4:
@@
$+\lambda_{power}\cdot\max\left\{0,\sum_{t=1}^{T}s_{\mathrm{p}}(t)^{2}-\gamma\right\}+\lambda_{SDF}\mathcal{L}_{SDF}(\mathbf{x}_{\mathrm{p}})$
$s_{\mathrm{p}}$ is the perturbation source-sound attained from Equation 3, $r(t;\mathbf{x}_{\mathrm{p}},\mathbf{x}_{\mu})$ is the RIR between sound-source location $\mathbf{x}_{\mathrm{p}}$ and microphone location $\mathbf{x}_{\mu}$, given by the pre-trained NAF, $\mathcal{F}$ is the pre-trained localization model with a set of microphones $\mathcal{M}$, where each microphone $\mu$ is located in location $\mathbf{x}_{\mu}$. $\mathbf{x}_{\mathrm{d}}$ represents actual location of the center of the drone, and $\beta,\gamma$ are the respective amplitude and power bounds from Section 3.3.2, for losses weighted by $\lambda_{amp}$ and $\lambda_{power}$. Location SDF is weighted by $\lambda_{SDF}$. While Equation 4 is formulated for non-targeted attacks, in Section 4.2 and in the supplementary we also explore targeted attacks, for which we optimize similarly to Equation 4, with the single modification of replacing $\mathbf{x}_{\mathrm{d}}$ with the designated target location.
@@ 3.3.4 Source Location Optimization
One potential difficulty in our formulation from Equation 4 is that there are cases where optimization of the perturbation source location $\mathbf{x}_{\mu}$ within the scene may be difficult or altogether impossible. The reason is that optimizing $\mathbf{x}_{\mu}$ necessitates differentiation through the acoustic model of the environment (namely, in our case, the NAF). To address cases where such a differentiable model is not available or where the added computational costs incurred by querying and backpropagating through such, usually large, models are too heavy (see supplementary for time and memory footprint in our experiments), in our experiments, we also analyze universal attacks with a fixed perturbation source location in the center of the room, only optimizing for the perturbation sound. In this case, the perturbation impulse response for every microphone location is constant, circumventing repeated forward and backward evaluation of the NAF. As Section 4.2 shows, we observe only a marginal decrease in attack prowess when waiving source-location optimization. We attribute this outcome to the fact that the source location should not drastically change the set of attainable location-wise perturbations potentially sampled at the sensors, but only the specific perturbative source-emitted waveform able to induce them over a given scene location. Our adversarial pipeline is illustrated in Figure 2.
@@ 3.4 Phase Modulation Perturbation Delineation
In this section, we propose Phase Modulation Perturbation Delineation - a novel method for the utilization of phase modulation (Shor et al., 2024) for the recovery of the ”clean,” non-perturbed signal from the entire (presumably adversarially perturbed) acoustic sample as perceived by any of the drone’s sensors. Section 3.4.1 overviews the concept of phase modulation, originally developed in (Shor et al., 2024) for improving localization accuracy. In Section 3.4.2, we show how this mechanism can be further extended for perturbation and clean-signal separation.
@@ 3.4.1 Phase Modulation
In our acoustic localization setting, similar to that studied in (Shor et al., 2024), the drone’s propulsion system consists of a set of $\mathcal{R}$ rotors that rotate at constant and equal angular velocities. The resulting angular locations of each rotor through the duration of a single cycle spanning over $T$ seconds (namely, the azimuthal rotor shaft positions w.r.t some chosen starting point) can thus be described as a function $\varphi_{0}(t):[T]\rightarrow[0,2\pi]$ converting moments in time from each cycle to the corresponding angular location of rotor $i$.
The phase modulation mechanism from (Shor et al., 2024) proposes to actively optimize a per-rotor function of temporal offsets from $\varphi_{0}(t)$, termed a modulation function $\varphi_{i}(t):[T]\rightarrow[0,2\pi]$, for every rotor $i$. Every combination of such modulation functions alters the sound emitted by the propulsion system through time, and thus the rationale is to optimize the set of $\mathcal{R}$ such per-rotor modulation functions to generate input waveforms more useful for localization (under imposition of certain constraints over these modulation functions, ensuring kinematic feasibility and the drone’s stable flight).
@@ 3.4.2 Perturbation Delineation
We extend the concept of phase modulation (Section 3.4.1) to delineate the adversarial perturbation from the entire sampled input $s_{\mu}$ at a given microphone $\mu$. Let us denote the perturbation response of $s_{\mathrm{p}}$ at microphone $\mu$ by $\sigma_{\mathrm{p}}$. Upon sampling a perturbed waveform $s_{\mu}=s_{drone}+\sigma_{\mathrm{p}}$, the goal is to solve for $\sigma_{\mathrm{p}}$. Upon doing so, we can input the original non-perturbed signal $s_{drone}$ to the localization model, essentially nullifying the affect of the attack over localization performance.
Our strategy for achieving this would be to apply gradual temporal delays over $s_{drone}$ using the phase modulation mechanism, utilizing the fact the the perturbation sound remains constant under these modulations. Denote by $T$ the maximum of the period times of the drone and perturbation source. Since the drone cycle $T_{drone}$ must be an integer multiple of the perturbation source period $T_{pert}$ (Section 3.3.2), both $s_{drone}$ and $s_{\mathrm{p}}$ complete a periodic cycle (of one or more periods) within a duration of $T$ seconds.
@@
By performing this process for every value of $j\in[T]$, we can essentially recover $\sigma_{\mathrm{p}}(t)-\sigma_{\mathrm{p}}(t=0)$ for every value of $t$, giving us the perturbating waveform at the location of the microphone $\mu$. One degree of uncertainty which our algorithm does not allow solving for, is the value of $\sigma_{\mathrm{p}}$ at moment $t=0$. While we believe this value can be retrieved, for example by imposing correlation among several perturbation waveforms at different locations, we defer the study of this possibility to future work. For the scope of this work, as validated in Section 4.3, simple setting of $\sigma_{\mathrm{p}}(t=0)$ to be 00 everywhere in the environment reduces localization degradation almost completely. We stress that while our proposed algorithm is demonstrated on phase modulation of drone rotors, it is applicable for acoustic perturbation delineation of any agent capable of actively shaping the ”clean” sound, thus separating it from the constant perturbation.
@@
Figure 3 visually depicts our algorithm - a single drone period is illustrated under 3 different phase modulations - 0 (A), $\pi$ (B, corresponding to $\tau_{1}$ timesteps), and $\frac{pi}{2}$ (C, corresponding to $\tau_{2}$ timesteps). In each room the same perturbation sound (red) is observed, while for each modulation the same drone-emitted sound is observed up to a certain shift (blue). The superimposed sound sampled by the microphone (green) is shifted accordingly (w.r.t the 0-modulation room A) in each room. In the bottom part, we illustrate how the 0-modulation waveform is subtracted from each $\tau$-step modulated waveform to receive the original perturbation (red waveform) at timestep $\tau$, utilizing the fact the the perturbation is agnostic to phase modulation.
@@ 4.1 Experimental Setup
Following the experimental setting in (Luo et al., 2022), we pick a representative subset of acoustic environments from the Matterport3D (Chang et al., 2017) and Replica (Straub et al., 2019) datasets used in training of RIRs from Neural Acoustic Fields (NAFs)(Luo et al., 2022). In this section, we primarily present results from the apartment_2_frl environment (denoted apt, in this section). We also provide partial results for the office_4 and room_2 scenes in Figure 5, with the full set of results available in the supplementary material. These additional results exhibit trends consistent with those reported here. For each environment we train the clean localization model from (Shor et al., 2024), using the environment-fitted pre-trained NAF. In order to explore our developed approaches disjointly from any external, case-specific affects, in most experiments we assume zero interference from any acoustic signals, excluding the drone and the attacker. Nonetheless, in Section B.1 we also model the affect of white noise over adversarial success. Furthermore, our experiments we focus on the regressor from (Shor et al., 2024) since it is the only one completely reliant on the self-sound emitted by the drone’s propulsion system. Other methods (e.g. (Sun et al., 2023; He et al., 2023)) rely mostly on external speakers, and thus require the development of adversarial attacks accounting for both the speaker sound and propulsion sound simultaneously. We refer these direction of research to future work. Nonetheless, we emphasize that our developed attack and defense methods are applicable to any combination of drone configuration and purely-acoustic localization algorithm. Localization training details are stated in the supplementary.
@@ 4.2 Attack Results
Optimized attack RMS error results across different bounds are reported in Figure 4. Our proposed attack increases the mean localization error from slightly below 5% in the clean model, to 37.4% for highest amplitude and power bounds. Results also demonstrate saturation of attack efficacy with growth in the $\beta,\gamma$ bounds, where in both figures the largest and second-largest amplitude bounds $\beta=0.5,\beta=1$ intersect almost completely. This is also true for the power bounds, as all figures remain almost constant in transition from $\gamma=1$ to $\gamma=2$. This bounds the maximum efficacy of our PGD attack in bound of around $0.5$ for the amplitude and $\gamma=1$ for power. Lastly, corresponding to Section 3.3.4, we observe that results for location optimization (left) and fixed-source location experiments (right) are nearly identical. We deem the apparent agnostic nature of sound-source perturbation optimization to source location a notable conclusion for future related studies, especially given the increased computational costs (see supplementary material).
@@
Figure 4: Mean RMS with and without source location optimization - across varying amplitude and power bounds ($\beta\text{ and }\gamma$).
Figure 5: Clean, perturbed and delineation-recovered spatial error distribution for selected attack bounds - across different scenes (apt., room, and office), measured in mean RMS error across drone orientations. Optimized source location highlighted in blue.
@@
Figure 5 illustrates spatial RMS error distributions across varying scenes and $\beta,\gamma$ bounds for optimized source location perturbation optimization. We supply similar results for fixed-source location in the supplementary. ”Clean Scene” denotes non-perturbed localization, ”Attacked Scene” depicts localization subjected to the attack from Section 3.3 and ”Recovered Scene” shows localization with the perturbation delineation defense methods developed in Section 3.4. Results visually demonstrate the affect of our proposed attack over localization performance, spread across the majority of the environment in all three scenes. The smooth spatial distribution of error across each map implies that our universal adversarial attack method does not merely mislead the drone at isolated, specific locations but instead systematically disrupts the underlying localization capability. This consistency in error across diverse spatial coordinates implies that the attack successfully generalizes to attack the localization task as a whole, rather than exploiting isolated positional weaknesses. To address any concern of the similarity between fixed and optimized source in Figure 4 stemming from the perturbation source not changing in optimization, we stress that location-optimization experiments do demonstrate some change in source location under optimization. The restricted extent of this change from the center can be explained by the reduced affect of room geometry over perceived perturbation at the sensors as the source is closer the the sensor. This presumably pushed the perturbation source to remain close to the center during optimization, also insinuating that the conclusion from Figure 4 regarding the affect of source location optimization may change in highly non-convex rooms. We defer exploration of this possibility to future work.
@@
In our setting the exploration of targeted adversarial attacks is also of great importance, as such attacks would allow a potential attacker to direct the drone to a specific location within the environment. Our main analysis of the performance of our method in applying targeted attacks is brought in the supplementary, however in Figure 6 we present initial results of our developed attack’s potential in successful targeted attack. This potential is evident from the presented error heatmap in the attacked scene, depicting RMS error between prediction and adversarial target. For our largest-considered attack bounds, this error is almost zero all along the map. We refer the reader to the supplementary for further related results, including cases where our targeted attack has proven inefficient.
@@ 4.3 Defense Results
In this section we analyze the effectiveness of our proposed acoustic perturbation delineation algorithm from Section 3.4. Figure 5 shows qualitative support for our method’s capability in drastically reducing adversarial performance decay, producing recovered scenes almost indistinguishable from the corresponding clean scenes - for the most lenient attack constraints, our algorithm reduces localization error from 37% (Figure 4) to below 6%, only marginally higher than the error of 4.87% reported for clean localization. In the supplementary we also report the recovered RMS error statistics, similarly to Figure 4.
Figure 6: Targeted attack spatial distribution - of RMS error from desired adversarial target, for maximal attack bounds.
@@
We attribute the efficacy of our method to the fact that it is capable of retrieving the sensed perturbation with a only single scalar degree-of-freedom, that is the uncertainty in the sampled perturbation value at moment $t=0$. To further validate our method, in Section B.4 we provide thorough empirical analysis of this uncertainty, and conclude that this degree of uncertainty has marginal affect over post-reconstruction localization performance.
5 Discussion
5.1 Conclusion
- In this work we have presented an initial study of adversarial attacks within the task of acoustic localization of drones. In our study we first presented a formulation of adversarial attacks appropriate for the acoustic setting. We proceeded to benchmark and comprehensively analyze how universal acoustic perturbations, optimized according to our proposed formulation, affect localization performance under different conditions. Finally, we presented a novel algorithm for reliable recovery of the optimized adversarial perturbation, as sampled at the source. Results show that our formulated attack increases localization RMS error from, at most, 37%, and our recovery algorithm reduces this error back to slightly below 6% - only marginally higher than the clean error of 5%. While our discussion is mostly focused our drone localization, we deem it applicable for many autonomous-agent acoustic localization scenarios.
5.2 Limitations and Future Work
- While our work provides a foundational framework for adversarial attacks and defenses in acoustic localization, several limitations remain. First, in this study we restrict ourselves to the task of 2D acoustic localization. Many real-world localization scenarios incorporate a 6-DoF regression problem, calling for an extension of our contributions to higher-dimension localization. Furthermore, in order to conduct an impact-analysis focused on our developed attack, in this work we bound the discussion to explore a single adversarial sound-source in a mostly clean, controlled acoustic setting. We plan to explore the relaxation of those assumptions in the future. Lastly, while efficient, our perturbation delineation algorithm must be applied for an entire drone cycle for every moment in time $[0,T]$, which may be time-costly. We plan to explore expedited perturbation recovery in future work, as well as other adversarial defense algorithms.
Appendix A Supplementary Layout
In the following sections we report additional details complementing the main paper, ordered as follows - 1) Appendix B includes supplementary results for the three main types of experiments conducted in the paper -
We denote that our scatter-plot heatmap figures for the room_2 and office_4 environments may depict environment shapes slightly different than those known from (Chang et al., 2017; Straub et al., 2019) environments. The reason is we do not evaluate localization in locations where the drone cannot be present, due to it’s simulated physical dimensions from (Shor et al., 2024).
2) Appendix C provides in-depth implementation, optimization and training-regime details. 3) In Appendix D we discuss the added computational resources required in adversarial source location optimization, to further stress the importance of our findings from section 4.2, where we claim source-location optimization could prove cost-ineffective in some cases.
4) Appendix E further provide added details for our proposed perturbation delineation algorithm.
Appendix B Additional Results
B.1 Modeling Noise

### v2

@@ 3.3.3 Perturbation Propagation
Much like the "clean" sound emitted by the drone, our adversarial perturbation must also be propagated to the drone’s sensor array to be sampled by it. Our goal is to formulate a universal attack optimization scheme shaping the emitted signal at the source $s_{\mathrm{p}}$ so that the sampled response at the sensor $\mu$, denoted $\sigma_{\mathrm{p}}$, would optimally reduce localization accuracy. In our experiments, we consider both optimization of the emitted perturbation for a sound source located in the center of the room, as well as optimization of the emitted sound jointly with the perturbation speaker’s location $\mathbf{x}_{\mathrm{p}}\in\mathbb{R}^{2}$. For our localization quality measure, we choose Mean-Squared-Error (MSE), following the criterion from Serussi et al. (2024). Our universal attack optimization, in its most general form, is therefore formulated according to equation 4:
@@
$\hskip-42.67912pt+\lambda_{power}\cdot\max\left\{0,\sum_{t=1}^{T}s_{\mathrm{p}}(t)^{2}-\gamma\right\}+\lambda_{SDF}\mathcal{L}_{SDF}(\mathbf{x}_{\mathrm{p}})$
$s_{\mathrm{p}}$ is the perturbation source-sound attained from equation 3, $r(t;\mathbf{x}_{\mathrm{p}},\mathbf{x}_{\mu})$ is the RIR between sound-source location $\mathbf{x}_{\mathrm{p}}$ and microphone location $\mathbf{x}_{\mu}$, given by the pre-trained NAF, $\mathcal{F}$ is the pre-trained localization model with a set of microphones $\mathcal{M}$, where each microphone $\mu$ is located in location $\mathbf{x}_{\mu}$. $\mathbf{x}_{\mathrm{d}}$ represents actual location of the center of the drone, and $\beta,\gamma$ are the respective amplitude and power bounds from Section 3.3.2, for losses weighted by $\lambda_{amp}$ and $\lambda_{power}$. Location SDF is weighted by $\lambda_{SDF}$. While equation 4 is formulated for non-targeted attacks, in Section 4.2 and in Appendix B.1 we adapt our formulation to explore targeted attacks.
@@ 3.3.4 Source Location Optimization
One potential difficulty in our formulation from equation 4 is that there are cases where optimization of the perturbation source location $\mathbf{x}_{\mathrm{p}}$ within the scene may be difficult or altogether impossible. The reason is that optimizing $\mathbf{x}_{\mathrm{p}}$ necessitates differentiation through the acoustic model of the environment (namely, in our case, the NAF). To address cases where such a differentiable model is not available or where the added computational costs incurred by querying and backpropagating through such, usually large, models are too heavy (see Appendix D for time and memory footprint in our experiments), in our experiments, we also analyze universal attacks with a fixed perturbation source location in the center of the room. In this case we only optimize the perturbative sound. The alleviation in computational costs stems from the fact that the perturbation impulse response for every microphone location is constant, circumventing repeated forward and backward evaluation of the NAF model. As Section 4.2 shows, we observe only a marginal decrease in attack prowess when waiving source-location optimization. Our adversarial pipeline is illustrated in Figure 2.
@@ 3.4 Phase Modulation Perturbation Delineation
In this section, we propose Phase Modulation Perturbation Delineation - a novel method for the utilization of phase modulation (Serussi et al., 2024) for the recovery of the "clean," non-perturbed signal from the entire (presumably adversarially perturbed) acoustic sample as perceived by any of the drone’s sensors. Section 3.4.1 overviews the concept of phase modulation, originally used by Serussi et al. (2024) for improving localization accuracy. In Section 3.4.2, we show how this mechanism can be further extended for perturbation and clean-signal separation.
@@ 3.4.1 Phase Modulation
In our acoustic localization setting, similar to that studied in Serussi et al. (2024), the drone’s propulsion system consists of a set of $\mathcal{R}$ rotors that rotate at constant and equal angular velocities. The resulting angular locations of each rotor through the duration of a single cycle spanning over $T$ seconds (namely, the azimuthal rotor shaft positions w.r.t some chosen starting point) can thus be described as a function $\varphi_{0}(t):[T]\rightarrow[0,2\pi]$ converting moments in time from each cycle to the corresponding angular location of rotor $i$.
The phase modulation mechanism from Serussi et al. (2024) proposes to actively optimize a per-rotor function of temporal offsets from $\varphi_{0}(t)$, termed a modulation function $\varphi_{i}(t):[T]\rightarrow[0,2\pi]$, for every rotor $i$. Every combination of such modulation functions alters the sound emitted by the propulsion system through time, and thus the rationale is to optimize the set of $\mathcal{R}$ such per-rotor modulation functions to generate input waveforms more useful for localization (under imposition of certain constraints over these modulation functions, ensuring kinematic feasibility and the drone’s stable flight).
@@ 3.4.2 Perturbation Delineation
We extend the concept of phase modulation (Section 3.4.1) to delineate the adversarial perturbation from the entire sampled input $s_{\mu}$ at a given microphone $\mu$. Let us denote the perturbation response of $s_{\mathrm{p}}$ at microphone $\mu$ by $\sigma_{\mathrm{p}}$. Upon sampling a perturbed waveform $s_{\mu}=s_{drone}+\sigma_{\mathrm{p}}$, the goal is to solve for $\sigma_{\mathrm{p}}$. Upon doing so, we can input the original non-perturbed signal $s_{drone}$ to the localization model, essentially nullifying the effect of the attack over localization performance.
Our approach is applying gradual temporal delays over $s_{drone}$ using the phase modulation mechanism, utilizing the fact the perturbation sound remains constant under these modulations. Denote by $T$ the maximum of the period times of the drone and perturbation source. Since the drone cycle $T_{drone}$ must be an integer multiple of the perturbation source period $T_{pert}$ (Section 3.3.2), both $s_{drone}$ and $s_{\mathrm{p}}$ complete a periodic cycle (of one or more periods) within a duration of $T$ seconds.
@@
By performing this process for every value of $j\in[T]$, we can essentially recover $\sigma_{\mathrm{p}}(t)-\sigma_{\mathrm{p}}(t=0)$ for every value of $t$, giving us the perturbating waveform at the location of the microphone $\mu$. One degree of uncertainty which our algorithm does not allow solving for, is the value of $\sigma_{\mathrm{p}}$ at moment $t=0$. While we believe this value can be retrieved, for example by imposing correlation among several perturbation waveforms at different locations, we defer the study of this possibility to future work. For the scope of this work, as validated in Section 4.3, simple setting of $\sigma_{\mathrm{p}}(t=0)$ to be 0 everywhere in the environment reduces localization degradation almost completely. We stress that while our proposed algorithm is demonstrated on phase modulation of drone rotors, it is applicable for acoustic perturbation delineation of any agent capable of actively shaping the "clean" sound, thus separating it from the constant perturbation. This formulation is summarized in algorithm 1.
@@
Figure 3 visually depicts our algorithm - a single drone period is illustrated under 3 different phase modulations - 0 (A), $\pi$ (B, corresponding to $\tau_{1}$ timesteps), and $\frac{pi}{2}$ (C, corresponding to $\tau_{2}$ timesteps). In each room the same perturbation sound (red) is observed, while for each modulation the same drone-emitted sound is observed up to a certain shift (blue). The superimposed sound sampled by the microphone (green) is shifted accordingly (w.r.t the 0-modulation room A) in each room. In the bottom part, we illustrate how the 0-modulation waveform is subtracted from each $\tau$-step modulated waveform to receive the original perturbation (red waveform) at timestep $\tau$, utilizing the fact the perturbation is agnostic to phase modulation.
1:Perturbed waveform $s_{\mu}$, period time $T$
2:Recovered sampled perturbation waveform $s_{\mathrm{r}}$
3:Initialize $s_{\mathrm{r}}=[\ ]$ $\triangleright$ Recovered waveform
4:for $j=1$ to $T$ do
5:  sample $s_{\mu}(t;j)$ $\triangleright$ Modulate by $j$ timesteps
6:  update $s_{\mathrm{r}}(t=j)=s_{\mu}(t;j)-s_{\mu}(t;0)$ $\triangleright$ Perturbation at timestep $j$
7:end for
8:return $s_{\mathrm{r}}$
@@ 4.1 Experimental Setup
Following the experimental setting in Luo et al. (2022), we pick a representative subset of acoustic environments from the Matterport3D (Chang et al., 2017) and Replica (Straub et al., 2019) datasets used in training of RIRs from Neural Acoustic Fields (NAFs). In this section, we primarily present results from the apartment_2_frl environment (denoted apt, in this section). We also provide partial results for the office_4 and room_2 scenes in Figure 5, with the full set of results available in the supplementary material (Appendix B). For each environment we train the clean localization model from Serussi et al. (2024), using the environment-fitted pre-trained NAF. In order to explore our developed approaches disjointly from any external, case-specific affects, in most experiments we assume zero interference from any acoustic signals, excluding the drone and the attacker. Nonetheless, in Section 4.4 we also model the effect of white noise over adversarial success. Furthermore, in our experiments we focus on the localization model from Serussi et al. (2024) since it is the only one completely reliant on the self-sound emitted by the drone’s propulsion system. Other methods (e.g. Sun et al. (2023); He et al. (2023)) rely mostly on external speakers, and thus require the development of adversarial attacks accounting for both the speaker sound and propulsion sound simultaneously. We refer this direction of research to future work. Nonetheless, we emphasize that our developed attack and defense methods are applicable to any combination of drone configuration and acoustic localization algorithm. Localization training details are stated in Appendix C.
@@ 4.2 Attack Results
Optimized attack RMS error results across different bounds are reported in Figure 4. Our proposed attack increases the mean localization error from slightly below 5% in the clean model, to 37.4% for highest amplitude and power bounds. Results also demonstrate saturation of attack efficacy with growth in the $\beta,\gamma$ bounds, where in both figures the largest and second-largest amplitude bounds $\beta=0.5,\beta=1$ intersect almost completely. This is also true for the power bounds, as all figures remain almost constant in transition from $\gamma=1$ to $\gamma=2$. This bounds the maximum efficacy of our PGD attack in bound of around $0.5$ for the amplitude and $\gamma=1$ for power. Lastly, corresponding to Section 3.3.4, we observe that results for location optimization (left) and fixed-source location experiments (right) are nearly identical. We deem the apparent agnostic nature of sound-source perturbation optimization to source location a notable conclusion for future related studies, especially given the increased computational costs (see Appendix D).
@@
Figure 4: Mean RMS error with and without source location optimization across varying amplitude ($\beta$) and power ($\gamma$) bounds. Optimizing the perturbation source location yields negligible improvement over a fixed source in the room center.
Figure 5: Spatial error distribution (Mean RMS) for Clean (top), Attacked (middle), and Recovered (bottom) scenarios across three environments (apt., room, and office). The "Attacked" row demonstrates that the universal perturbation degrades localization performance uniformly across the entire environment (indicated by high RMS values), rather than exploiting specific positional weaknesses. The "Recovered" row shows that our phase-modulation defense successfully delineates the perturbation, restoring localization accuracy to near-clean levels. Blue markers indicate optimized source locations.
@@
Figure 5 illustrates spatial RMS error distributions across varying scenes and $\beta,\gamma$ bounds for optimized source location perturbation optimization. We supply similar results for fixed-source location in appendix B. "Clean Scene" denotes non-perturbed localization, "Attacked Scene" depicts localization subjected to the attack from Section 3.3 and "Recovered Scene" shows localization with the perturbation delineation defense methods developed in Section 3.4. Results visually demonstrate the effect of our proposed attack over localization performance, spread across the majority of the environment in all test scenes. The smooth spatial distribution of error across each map implies that our universal adversarial attack method manages to markedly diminish the underlying localization capability throughout the scene. This consistency in error across diverse spatial coordinates implies that the attack successfully generalizes to attack the localization task as a whole, rather than exploiting isolated positional weaknesses. To address any concern of the similarity between fixed and optimized source in Figure 4 stemming from the perturbation source not changing in optimization, we stress that location-optimization experiments do demonstrate some change in source location under optimization. The restricted extent of this change from the center can be explained by the reduced affect of room geometry over perceived perturbation at the sensors as the source is closer to the sensor. This pushes the perturbation source to remain close to the center during optimization, also insinuating that the conclusion from Figure 4 regarding the effect of source location optimization may change in highly non-convex rooms. We defer exploration of this possibility to future work.
@@
In our setting the exploration of targeted adversarial attacks is also of great importance, as such attacks would allow a potential attacker to direct the drone to a specific location within the environment. Our main analysis of the performance of our method in applying targeted attacks is brought in the supplementary (Appendix B.1), however in Figure 6 we present initial results of our developed attack’s potential in successful targeted attack. This potential is evident from the presented error heatmap in the attacked scene, depicting RMS error between prediction and adversarial target. For our largest-considered attack bounds, this error is almost zero all along the map. We refer the reader to the supplementary (Appendix B) for further related results, including cases where our targeted attack has proven inefficient.
@@ 4.3 Defense Results
In this section we analyze the effectiveness of our proposed acoustic perturbation delineation algorithm from Section 3.4. Figure 5 shows qualitative support for our method’s capability in drastically reducing adversarial performance decay, producing recovered scenes almost indistinguishable from the corresponding clean scenes - for the most lenient attack constraints, our algorithm reduces localization error from 37% (Figure 4) to below 6%, only marginally higher than the error of 4.87% reported for clean localization. In Appendix E we also report the recovered RMS error statistics, similarly to Figure 4.
Figure 6: Targeted attack spatial distribution - of RMS error relative to the adversarial target (indicated by the blue dot) under a targeted attack with maximal bounds. The uniformly low error (black regions) indicates a highly successful attack, where the drone is consistently deceived into localizing itself at the adversary’s chosen coordinates regardless of its actual position in the room.
@@
We attribute the efficacy of our method to the fact that it is capable of retrieving the sensed perturbation with a only single scalar degree-of-freedom, that is the uncertainty in the sampled perturbation value at moment $t=0$. To further validate our method, in the appendix we provide thorough empirical analysis of this uncertainty, and conclude that this degree of uncertainty has marginal affect over post-reconstruction localization performance.
4.4 Noise Modeling